OpenStudy (anonymous):

Find the derivative of the function. g(x)=intergral from 2x to 3x of (u^(2)-1)/(u^(2)+1)du

OpenStudy (anonymous):

it gives the hint of \[\int\limits_{2x}^{3x}f(u)du=\int\limits_{2x}^{0}f(u)du+\int\limits_{0}^{3x}f(u)du\]

OpenStudy (anonymous):

the problem is also \[g(x)=\int\limits_{2x}^{3x}\frac{ u^2-1 }{ u^2+1}\]